OpenStudy (bigbosssaint21):

simplify the expression as much as possible and enter the result below ((-125)^3)^1/9

OpenStudy (welshfella):

use the rule ( x^n)^m = x^nm

OpenStudy (welshfella):

so we have (-125)^(3 * 1/9) = (-125)^(1/3) can you continue?